OpenStudy (kayluh123):

You need a 45% alcohol solution. On hand, you have a 125 mL of a 10% alcohol mixture. You also have 70% alcohol mixture. How much of the 70% mixture will you need to add to obtain the desired solution?

OpenStudy (aaronq):

So if the amount of solute is given by multiplying the concentration by the volume, then we can use the two solutions to give the total, like so, \(\sf solute~ 1+solute~2=total ~solute\) \(\sf C_1*V_1+C_2*V_2=C_T*V_T\) also, \(V_T=V_1+V_2\) \(\sf C_1*V_1+C_2*V_2=C_T*(V_1+V_2) \) plug in your values: \(\sf (0.1*0.125~L)+0.7*V_2=0.45*(0.125~L+V_2) \) \(\sf V_2=0.175 ~L =175~mL \) so add 175 mL of the 70% solution

OpenStudy (kayluh123):

thank you!!!

OpenStudy (aaronq):

no problem !